# DDGS-CT: Direction-Disentangled Gaussian Splatting for Realistic Volume Rendering

### v1

@@ 4.1 Experimental Protocol
For evaluation, we set the decomposition degree to $L$=1, the feature dimension to $k$=8, and initial cloud sizes to $n_{1}$=15,000 and $n_{2}$=10,000. We apply a loss weight $\lambda=0.2$. We choose the default threshold level (i.e., the average of the minimum and maximum volume) for the marching-cubes algorithm. DDGS training is performed on a single NVIDIA A100 using an Adam optimizer Kingma and Ba (2014) with a learning rate of $1.25\times 10^{-4}$ for $\bm{b}^{\text{iso}}$ and $\bm{B}^{\text{dir}}$ and $2.5\times 10^{-3}$ for $\bm{f}^{\text{iso}}$ and $\bm{f}^{\text{dir}}$. Default 3DGS Kerbl et al. (2023) learning rates are applied to the remaining parameters.
@@
We consider 4 datasets. (1) NAF-CT Zha et al. (2022) includes four CT images of abdomen, chest, foot, and jaw. For each, we adopt TIGRE Biguri et al. (2016) to sample 50 evenly-distributed projections for training and 50 randomly-distributed ones for testing, in the range of [90°- 90  °, 90°90  °] (scatter-free DRRs). (2) CTPelvic1K Liu et al. (2021) is a large dataset of pelvic CT images. We consider the first 10 scans of the sub-dataset6 in our experiments. We adopt DeepDRR Unberath et al. (2018) to generate 60 training (evenly-distributed) and 60 (randomly-distributed) testing DRRs (in the range of [60°- 60  °, 60°60  °]), taking advantage of the scattering-modeling capability of DeepDRR to render realistic X-ray data. (3) Ljubljana Pernus et al. (2013) is a clinical dataset of 10 patients undergoing neurovascular surgery. Each patient underwent one CT and two X-ray angiography scans. The pose of each X-ray is also provided. Following DiffPose strategy Gopalakrishnan et al. (2023), we randomly sample 900 projections for training and 100 for testing. (4) Finally, we provide some additional qualitative results on DeepFluoro Grupp et al. (2020), a collection of pelvic X-rays and CT images from six cadavers.
@@
We evaluate our method in two settings: novel-view synthesis and intraoperative 2D/3D image registration (i.e., pose estimation). For novel-view synthesis, we use the standard PSNR and SSIM as metrics. For intraoperative 2D/3D image registration, we consider the error both in terms of rotation (angular distance) and translation (Euclidean distance). We further measure the clinically-relevant registration accuracy of key anatomical landmarks for Ljubljana scans Pernus et al. (2013) where they are provided. Following Gopalakrishnan et al. (2023), we compute the target registration error (TRE) w.r.t. the positioning of these landmarks after registration.
@@ 4.2 Novel-View Synthesis
3DGS ($L=1$) X-Gaussian ($k=32$) DDGS (Ours, $L=1,k=8$)
@@
We compare to 3DGS Kerbl et al. (2023) and X-Gaussian Cai et al. (2024) on the NAF-CT Zha et al. (2022) and CTPelvic1K Liu et al. (2021) datasets for the novel-view synthesis. For each method, we adopt the hyper-parameters recommended by their respective authors (e.g., $L=1$ and 333 for normal 3DGS Kerbl et al. (2023) and $k=8$ and 323232 for X-Gaussian Cai et al. (2024)).
@@
As shown in Table 1, for NAF-CT Zha et al. (2022) (without scatter-related noise), our DDGS outperforms both 3DGS and X-Gaussian in terms of PSNR and SSIM. For CTPelvic1K Liu et al. (2021) (with scatter), our DDGS significantly outperforms 3DGS and X-Gaussian in both PSNR and SSIM while using considerably fewer points. Therefore, our DDGS method is more effective in simulating realistic X-ray images from CT scans.
Furthermore, we compare our DDGS (with degrees $L=1$ and $L=3$ and feature dimension $k=8$) with traditional 3DGS (degrees $L=1$ and $L=3$) and X-Gaussian (feature dimensions $k=8$ and $k=32$) on the 001 and 002 data of CTPelvic1K Liu et al. (2021) at iterations of 500, 2000, 7000, 15,000, and 30,000, as shown in Table 2. Our model with $L=1$ and $k=8$ performs comparably to 3DGS with $L=3$ and outperforms X-Gaussian with $k=32$ after 15,000 iterations. Our model with $L=3$ achieves the best performance overall. Figure 3 presents a qualitative comparison of the testing set, illustrating both the synthetic views and the generated 3D Gaussian points.
@@
Table 4: Impact of initialization strategies on final image quality, measured CTPelvic1K scans.
001 002 003 004 005 006 007 008 009 010 avg
Radiodensity-based (Ours) 37.88 38.43 38.28 40.35 40.36 38.26 38.87 39.81 39.87 38.73 39.08
Random (uniform dis.) 37.17 38.36 38.15 40.09 39.97 38.36 38.69 39.37 39.85 38.84 38.89
Even sampling 37.33 38.51 37.83 39.58 39.12 38.27 38.87 39.88 39.46 38.95 38.78
@@
Table 6: Accuracy of image registration on the CTPelvic1K Liu et al. (2021) dataset.
3DGS X-Gaussian ($k$=8) X-Gaussian ($k$=32) DDGS (Ours)
@@
Table 3 demonstrates the effectiveness of our direction-disentangled representation for 3D Gaussians. For several settings, we observe performance degradation when learning the direction-dependent and direction-independent components in the same 3D Gaussians (direct-entangled), learning the direction-independent 3D Gaussians alone (direct-independent), or learning the direction-dependent 3D Gaussians alone (direct-dependent). Additionally, we decouple the direction-dependent and direction-independent components in 3DGS Kerbl et al. (2023) (3DGS-disentangled), which also showed lower performance compared to our learnable features. Figure 4 shows the rendered views from disentangled 3D Gaussians, highlighting that the direction-dependent 3D Gaussians focus more on bones or anatomical structures, while the direction-independent 3D Gaussians focus more on the background.
Table 4 shows the effectiveness of our proposed initialization scheme (radiodensity-aware) compared to other initialization methods. Our scheme outperforms both the uniformly random sampling used in 3DGS Kerbl et al. (2023) and the even sampling method used in X-Gaussian Cai et al. (2024).
@@ 4.3 Application to Downstream Task (2D/3D CT Image Registration)
We choose the first five testing images on the 001, 002, and 003 CT data of CTPelvic1K Liu et al. (2021) for image registration. Considering this task as an optimization-based problem that can be solved via inverse graphics (i.e., rendering synthetic images based on predicted poses and comparing to the target image, then backpropagating the difference to improve the pose prediction), we adopt the framework from iComMA Sun et al. (2023), replacing their image rendered by DDGS. We use Adam optimizer with a learning rate of 0.05. We use the isocenter pose as the initialization pose. As shown in Table 6 and Figure 5, our DDGS achieves the lowest rotation and translation errors.
Finally, we also consider the experimental protocol proposed in DiffPose Gopalakrishnan et al. (2023), where the authors rely on a pretrained pose-regression CNN to get a first rough estimation of the scanner pose, before refining said pose via a gradient-descent based optimization, leveraging their differentiable DRR renderer Gopalakrishnan and Golland (2022). We adopt their pose-initialization network and testbed for the Ljubljana dataset, replacing their DiffDRR analytical renderer Gopalakrishnan and Golland (2022) by our DDGS, and compare to them in terms of runtime performance and landmark-registration accuracy. The results, reported in Table 7, highlight the convergence boost brought by our lighter solution.
@@ 5 Conclusion
Limitations. It should be noted that, similar to previous methods, our model expects an offline DRR renderer to produce realistic ground truth to train DDGS, which can be a limitation for some scenarios (e.g., when such renderer or certain parameters required for realistic rendering are not available). Our direction-dependent function relating to anisotropic X-ray effects may also fail to approximate complex multi-bounce light scenarios. Recent developments in $N$-dimensional Gaussian splatting Diolatzis et al. (2024) show promising potential for modeling such cases.
@@
Societal Impact. The above limitation may restrict the adoption of DDGS to specific use cases, and our method should undergo a clinical evaluation of the generated images in terms of anatomical accuracy. Nonetheless, we believe that our proposed method can positively impact both the computer vision and medical communities by providing a more efficient and versatile DRR tool.

### v2

@@ 4.1 Experimental Protocol
For evaluation, we set the decomposition degree to $L$=1, the feature dimension to $k$=8, and initial cloud sizes to $n_{1}$=15,000 and $n_{2}$=10,000. We apply a loss weight $\lambda=0.2$. We choose the default threshold level (i.e., the average of the minimum and maximum volume) for the marching-cubes algorithm. DDGS training is performed on a single NVIDIA A100 using an Adam optimizer [18] with a learning rate of $1.25\times 10^{-4}$ for $\bm{b}^{\text{iso}}$ and $\bm{B}^{\text{dir}}$ and $2.5\times 10^{-3}$ for $\bm{f}^{\text{iso}}$ and $\bm{f}^{\text{dir}}$. Default 3DGS [17] learning rates are applied to the remaining parameters.
@@
We consider 4 datasets. (1) NAF-CT [42] includes four CT images of abdomen, chest, foot, and jaw. For each, we adopt TIGRE [5] to sample 50 evenly-distributed projections for training and 50 randomly-distributed ones for testing, in the range of [90°-90°- 90  °, 90°90°90  °] (scatter-free DRRs). (2) CTPelvic1K [19] is a large dataset of pelvic CT images. We consider the first 10 scans of the sub-dataset6 in our experiments. We adopt DeepDRR [35] to generate 60 training (evenly-distributed) and 60 (randomly-distributed) testing DRRs (in the range of [60°-60°- 60  °, 60°60°60  °]), taking advantage of the scattering-modeling capability of DeepDRR to render realistic X-ray data. (3) Ljubljana [26] is a clinical dataset of 10 patients undergoing neurovascular surgery. Each patient underwent one CT and two X-ray angiography scans. The pose of each X-ray is also provided. Following DiffPose strategy [12], we randomly sample 900 projections for training and 100 for testing. (4) Finally, we provide some additional qualitative results on DeepFluoro [13], a collection of pelvic X-rays and CT images from six cadavers.
@@
We evaluate our method in two settings: novel-view synthesis and intraoperative 2D/3D image registration (i.e., pose estimation). For novel-view synthesis, we use the standard PSNR and SSIM as metrics. For intraoperative 2D/3D image registration, we consider the error both in terms of rotation (angular distance) and translation (Euclidean distance). We further measure the clinically-relevant registration accuracy of key anatomical landmarks for Ljubljana scans [26] where they are provided. Following [12], we compute the target registration error (TRE) w.r.t. the positioning of these landmarks after registration.
@@ 4.2 Novel-View Synthesis
3DGS [17] ($L=1$) X-Gaussian [7] ($k=32$) DDGS (Ours, $L=1,k=8$)
@@
We compare to 3DGS [17] and X-Gaussian [7] on the NAF-CT [42] and CTPelvic1K [19] datasets for the novel-view synthesis. For each method, we adopt the hyper-parameters recommended by their respective authors (e.g., $L=1$ and $3$ for normal 3DGS [17] and $k=8$ and $32$ for X-Gaussian [7]).
@@
As shown in Table 1, for NAF-CT [42] (without scatter-related noise), our DDGS outperforms both 3DGS and X-Gaussian in terms of PSNR and SSIM. For CTPelvic1K [19] (with scatter), our DDGS significantly outperforms 3DGS and X-Gaussian in both PSNR and SSIM while using considerably fewer points. Therefore, our DDGS method is more effective in simulating realistic X-ray images from CT scans.
Furthermore, we compare our DDGS (with degrees $L=1$ and $L=3$ and feature dimension $k=8$) with traditional 3DGS (degrees $L=1$ and $L=3$) and X-Gaussian (feature dimensions $k=8$ and $k=32$) on the 001 and 002 data of CTPelvic1K [19] at iterations of 500, 2000, 7000, 15,000, and 30,000, as shown in Table 2. Our model with $L=1$ and $k=8$ performs comparably to 3DGS with $L=3$ and outperforms X-Gaussian with $k=32$ after 15,000 iterations. Experiments on more data can be found in Table S1. Our model with $L=3$ achieves the best performance overall. Figure 3 presents a qualitative comparison of the testing set, illustrating both the synthetic views and the generated 3D Gaussian points.
It should also be highlighted that, even though both the original CT scan (as voxel grid) and our 3DGS-based solution are explicit representations, the latter is significantly more compact. E.g., our model can represent a CTPelvic1K scan with only 42,625 Gaussians (c.f. Table 1), defined by 19 float values each (3D position/rotation + 3D covariance + opacity + feature vector) and a shared 32-dimensional basis vector $\mathbf{b}$, so 809,907 float values in total; whereas the original CT scans are each composed of $512\times 512\times 500=93,363,200$ values. Indeed, the voxel data may contain large homogeneous regions, which can be approximated with few Gaussians, hence a significant compression rate.
Finally, similar to [12], we qualitatively compare the results of recent DRR methods with actual X-ray images, for datasets providing both CT scans and real, posed projections (DeepFluoro [13], Ljubljana [26]). Results are shared in Figure 4, with our method achieving state-of-the-art PSNR. Further analyses are provided in Appendix B.
@@
Figure 4: Qualitative comparison of DRRs and real scans from DeepFluoro and Ljubljana datasets.
@@
Table 4: Impact of initialization strategies on image quality during training, i.e., measured on CTPelvic1K scans w.r.t. training iterations.
Iteration Strategy 001 002 003 004 005 006 007 008 009 010 avg
500 RADS (ours) 25.86 27.72 25.24 28.52 30.52 28.31 29.23 27.92 28.95 29.20 28.15
Random 25.29 26.73 24.66 27.40 28.46 26.35 27.23 26.43 27.34 28.53 26.84
Even 24.96 26.64 24.40 26.60 27.54 26.63 25.06 26.01 26.15 26.92 26.09
7,000 RADS (ours) 35.02 37.06 36.47 36.49 36.79 34.52 35.56 36.03 36.44 35.55 35.99
Random 34.39 35.78 34.18 35.83 36.37 34.36 35.20 35.74 36.30 35.26 35.34
Even 34.22 36.07 33.65 36.31 36.64 34.13 34.97 35.85 35.61 35.33 35.28
30,000 RADS (ours) 37.88 38.43 38.28 40.35 40.36 38.26 38.87 39.81 39.87 38.73 39.08
Random 37.17 38.36 38.15 40.09 39.97 38.36 38.69 39.37 39.85 38.84 38.89
Even 37.33 38.51 37.83 39.58 39.12 38.27 38.87 39.88 39.46 38.95 38.78
@@
Table 6: Accuracy of image registration on the CTPelvic1K [19] dataset.
3DGS [17] X-Gaussian [7] ($k$=8) X-Gaussian [7] ($k$=32) DDGS (Ours)
@@
Table 3 demonstrates the effectiveness of our direction-disentangled representation for 3D Gaussians. For several settings, we observe performance degradation when learning the direction-dependent and direction-independent components in the same 3D Gaussians (direct-entangled), learning the direction-independent 3D Gaussians alone (direct-independent), or learning the direction-dependent 3D Gaussians alone (direct-dependent). Additionally, we decouple the direction-dependent and direction-independent components in 3DGS [17] (3DGS-disentangled), which also showed lower performance compared to our learnable features. Figure 5 shows the rendered views from disentangled 3D Gaussians, highlighting that the direction-dependent 3D Gaussians focus more on bones or anatomical structures, while the direction-independent 3D Gaussians focus more on the background.
Table 4 shows the effectiveness of our proposed initialization scheme (radiodensity-aware) compared to other initialization methods. Our scheme outperforms both the uniformly random sampling used in 3DGS [17] and the even sampling method used in X-Gaussian [7].
@@ 4.3 Application to Downstream Task (2D/3D CT Image Registration)
We choose the first five testing images on the 001, 002, and 003 CT data of CTPelvic1K [19] for image registration. Considering this task as an optimization-based problem that can be solved via inverse graphics (i.e., rendering synthetic images based on predicted poses and comparing to the target image, then backpropagating the difference to improve the pose prediction), we adopt the framework from iComMA [32], replacing their image rendered by DDGS. We use Adam optimizer with a learning rate of 0.05. We use the isocenter pose as the initialization pose. As shown in Table 6 and Figure 6, our DDGS achieves the lowest rotation and translation errors.
Finally, we also consider the experimental protocol proposed in DiffPose [12], where the authors rely on a pretrained pose-regression CNN to get a first rough estimation of the scanner pose, before refining said pose via a gradient-descent based optimization, leveraging their differentiable DRR renderer [11]. We adopt their pose-initialization network and testbed for the Ljubljana dataset, replacing their DiffDRR analytical renderer [11] by our DDGS, and compare to them in terms of runtime performance and landmark-registration accuracy. The results, reported in Table 7, highlight the convergence boost brought by our lighter solution.
@@ 5 Conclusion
Limitations. It should be noted that, similar to previous methods, our model expects an offline DRR renderer to produce realistic ground truth to train DDGS, which can be a limitation for some scenarios (e.g., when such renderer or certain parameters required for realistic rendering are not available). Our direction-dependent function relating to anisotropic X-ray effects may also fail to approximate complex multi-bounce light scenarios. Recent developments in $N$-dimensional Gaussian splatting [8] show promising potential for modeling such cases. We should also note that the quality of DRRs is bound to the quality of the input CT scans. This issue affects both DRR methods and Monte-Carlo simulations (as the latter relies on the segmentation of CT scans into materials-specific regions – if a 3D scan is noisy, so will the material segmentation and simulation results). When our method uses simulators to get target images, it will face similar noise issues, though, our experiments show that DDGS barely introduces additional noise. Compensating for CT noise is an interesting, under-explored research direction that could benefit all DRR methods; but, we believe it is beyond the scope of our study.
@@
Societal Impact. The above limitations may restrict the adoption of DDGS to specific use cases, and our method should undergo a clinical evaluation of the generated images in terms of anatomical accuracy. Nonetheless, we believe that our proposed method can positively impact both the computer vision and medical communities by providing a more efficient and versatile DRR tool.
Supplementary Material
In this supplementary material, we provide further methodological context and showcase additional quantitative and qualitative results to highlight further the contributions claimed in the paper.
Appendix A Methodological Insight
A.1 Transmittance in Natural vs. X-Ray Imaging
The exponential transmittance model $T(t)$ used in NeRF and Gaussian-splatting (GS) solutions for natural scenes—to describe the light attenuation as it travels through a medium from point $r(t_{0})$ to $r(t)$—is based on Beer-Lambert law [16, 23, 37, 33]:
$T(t)=\exp(-\int_{t_{0}}^{t}a(r(s))\delta s),$ (4)
where $r(t)=o+td$ is a 3D point at step $t$ on a ray of origin $o$ and direction $d$, and $a(p)$ is the medium’s absorption/extinction (linked to its density) at point $p$. Assuming a piece-wise homogeneous medium (with $a_{i}$ the absorption of homogeneous segment $l$ of length $\delta_{l}$), this transmittance can be rewritten:
$T(t)=\exp(-\sum_{l=1}^{l_{t}}a_{l}\delta_{l})=\prod_{l=1}^{l_{t}}{(1-\sigma_{l})},$ (5)
as in our Equation 1; with $\sigma_{l}=1-\exp(-a_{l}\delta_{n})$ and $l_{t}$ the end segment containing $t$. See [33] for proof.
X-ray attenuation from photoelectric effect (PE) follows the same model: at each step through the medium (e.g., voxel), the probability of photon extinction (i.e., averaged attenuation) is proportional to the atomic density. E.g., if a ray travels through only 2 voxels, and if it is attenuated by $a_{1}=50$% at step 1 then $a_{2}=50$% again at step 2, then its total attenuation is 75% (c.f. $(100-(100-50)^{2})$%), in accordance to Beer-Lambert.
Hence, Equation 1 (3DGS rendering) could be directly applied to DDR synthesis (assuming standard neglog scaling of absorption values into pixel ones [12, 35]), if we were to adopt a simplified model of X-ray imaging which only considers isotropic absorption (i.e., following the Beer-Lambert law) and optionally omit the term $c_{j}\in\mathbb{R}^{K}$ (view-dependent radiance). This is the simplified model proposed in prior work [7].
In our paper, we propose to go beyond the isotropic simplification in X-Gaussian [7] and further account for anisotropic scattering of X-rays. This is why we preserve $c_{j}$ and decompose it into two terms: an isotropic term $c^{\text{iso}}$ and direction-dependent non-linear term $c^{\text{dir}}$ (Equation 3). Note that we also fix the dimension $K$ of these variables (i.e., the number of output channels) to 1 (monochromatic DRR) instead of three in natural 3DGS (RGB).
A.2 Joint Gaussian Rasterization
As explained in Section 3, we define two distinct functions for the absorption contribution of the two Gaussian sets: one function is isotropic to approximate average radio-absorption ($c_{i}^{\text{iso}}$), and one function is direction-dependent to approximate the contribution of anisotropic Compton scattering to the image ($c_{j}^{\text{dir}}$), c.f. Equation 3.
These two distinct Gaussian sets could be rasterized separately, as done in Figure 5 as qualitative ablation. However, the two sets need to be rendered together to ensure correct simulation. Otherwise, Gaussians from one set occluded by others belonging to the other set could incorrectly contribute to the ray absorption. We refer the readers to Figure S1 for an explanatory diagram. There, point $g_{2}^{\text{dir}}$ should not contribute to the final pixel value, as it is occluded by $g_{2}^{\text{iso}}$ (i.e., it absorbs all the remaining energy before the ray can reach $g_{2}^{\text{dir}}$). This could not be properly simulated if each Gaussian set is rasterized separately.
By having 2 sets of Gaussians with distinct yet complementary contribution functions (isotropic versus direction-dependent functions), our solution can better model DRR imaging. During optimization, each Gaussian set will approximate its respective imaging effect, conditioned by its respective function.
We randomly split $\mathbf{P}^{\text{mc}}$ (3D points sampled based on the CT scan distribution) into 2 sets, each subset contributing to the initialization positions of $g_{i}^{\text{iso}}$ and $g_{j}^{\text{dir}}$ (half the 3D points in $\mathbf{P}^{\text{mc}}$ will serve as initialization for $g_{i}^{\text{iso}}$, along with the entire $\mathbf{P}^{\text{mc}}$ set; and the other half of $\mathbf{P}^{\text{mc}}$ will serve as initialization for $g_{j}^{\text{dir}}$). After initialization, during 3DGS-based model optimization, the points in each set of Gaussians can move/split/drop independently (along with the optimization of their respective covariance/opacity/feature values). Since the optimization of each set (isotropic and anisotropic) is conditioned on their respective differentiable absorption function, they will acquire distinct properties.
Figure S1: Importance of joint rasterization of Gaussian sets (simplified diagram).
Appendix B Additional Experiments
B.1 Additional Comparison and Ablation Results
To further support the results in Table 2 and Table 3, we evaluate our method on all the 10 selected data from the CTPelvic1K dataset, as shown in Table S1.
Table S1: Novel-view synthesis evaluation on CTPelvic1K data.
DDGS (ours) ($L=1$, $k=8$) 3DGS ($L=1$) X-Gaussian ($k=8$) DDGS (ours) ($L=3$, $k=8$) 3DGS ($L=3$) X-Gaussian ($k=32$) direct-entangled c.f. Ablation Study
PNSR SSIM PNSR SSIM PNSR SSIM PNSR SSIM PNSR SSIM PNSR SSIM PNSR SSIM
001 38.04 .984 35.36 .970 33.39 .947 38.15 .985 37.12 .980 36.83 .979 36.17 .977
002 38.30 .987 37.25 .982 36.04 .978 38.97 .989 38.75 .988 37.89 .986 37.32 .984
003 38.21 .983 35.81 .974 33.99 .965 39.42 .986 39.09 .984 35.84 .976 35.32 .975
004 40.54 .987 38.43 .980 37.18 .974 40.74 .987 40.69 .986 39.06 .984 39.70 .985
005 39.88 .987 39.44 .985 37.50 .978 40.02 .987 40.68 .989 39.24 .985 39.36 .986
006 38.12 .986 36.69 .982 35.03 .977 38.79 .986 38.65 .987 37.14 .983 36.89 .983
007 38.91 .987 36.88 .978 35.51 .970 39.37 .988 38.58 .984 38.10 .987 37.72 .985
008 39.83 .982 37.85 .979 36.70 .973 40.19 .986 39.70 .986 38.57 .982 38.53 .982
009 39.67 .987 37.94 .980 37.60 .979 40.35 .988 39.16 .983 38.99 .987 38.74 .985
010 38.60 .987 37.50 .982 36.30 .978 39.18 .988 39.13 .986 37.89 .986 37.20 .985
avg 39.01 .986 37.32 .979 35.92 .972 39.52 .987 39.16 .985 37.96 .984 37.70 .983
B.2 Comparison to Real X-ray Projections
The main purpose of DRRs is to provide quick visualization to clinicians (with the key metrics being speed and visibility), as well as to integrate larger imaging applications (with priority given again to speed and feature-level similarity w.r.t. real data). Therefore, prior DRR papers (traditional [11, 30, 31] or GS-based [7, 25]) mostly evaluate on downstream tasks (e.g., pose registration). A few evaluate the image quality compared to other DRR tools (DiffDRR [11] compares to Siddon’s and Plastimatch [30]; [7] to other NVS baselines). We found that only DiffPose (based on DiffDRR) [12] provides a qualitative comparison to some real scans.
Quantitative comparison to real data is challenging. Not all imaging parameters are usually provided or accurate enough to create matching DRRs (w.r.t. intensity, CT pose, etc.). It is especially hard to render DRRs aligned with real scans with existing Monte-Carlo (MC) simulation tools [2, 3, 4, 15]; which is likely why no DRR paper has performed such evaluation. Their interfaces and custom coordinate conventions are not compatible with pose annotations in public CT/X-ray datasets, and their documentation/support is lacking. Bridging the convention gap between MC and analytical communities would greatly benefit our domain and could be the focus of our next effort.
Nevertheless, we do provide an indirect comparison to real projections. E.g., the registration experiments (Table 6) implicitly provide insight into the similarity between real and synthetic data, as pose optimization is done by comparing DRRs to real target scans.
In this supplementary, we share a more direct comparison. We use our DDGS-based registration method to refine GT poses in Ljubljana data (c.f. aforementioned GT inaccuracies), then use the refined poses to render and compare DRRs to real scans. Results can be found in Table S2 and Figure 4. E.g., when zooming into the images of the latter figure, one can observe that DiffDRR suffers from pixelization, unlike GS methods.
Table S2: Quantitative comparison of analytical DRRs to real images from Ljubljana dataset.
Methods SSIM PSNR
mean med std mean med std
DiffDRR [11] .404 .440 .179 24.16 24.35 2.82
3DGS [17] .409 .518 .199 22.57 23.44 2.78
X-Gaussian [7] .409 .521 .198 22.58 23.42 2.78
DDGS (ours) .459 .558 .204 25.61 27.35 3.48